OpenStudy (goldeverything):

At a bake sale, Jake raised one hundred one dollars and fifty cents toward his trip to soccer camp. When combined with another fundraiser, he has raised a total of two hundred dollars and twenty-five cents toward his trip. How much did Jake raise in the other fundraiser? ninety-eight dollars and seventy-five cents one hundred one dollars and fifty cents two hundred dollars and twenty-five cents three hundred one dollars and seventy-five cents MEDAL WILL BE AWARDED

OpenStudy (anonymous):

$101.50 + x dollars from the other fundraiser = $200.25

OpenStudy (anonymous):

So... 101.5 + x = 200.25 then... x = 200.25-101.5 = ...